OpenStudy (anonymous):

Three-fourths of a number increased by one-half is equal to fourteen. Find the number. 10/3/8 18 10/1/8 19/1/3

OpenStudy (misssunshinexxoxo):

What do you believe the answer is?

OpenStudy (anonymous):

18?

OpenStudy (anonymous):

yes

OpenStudy (misssunshinexxoxo):

That is correct; 18 is the solution

OpenStudy (texaschic101):

3/4n + 1/2 = 14 3/4n = 14 - 1/2 3/4n = 28/2 - 1/2 3/4n = 27/2 n = (27/2) / (3/4) n = 27/2 * 4/3 n = 108/6 n = 18